OpenStudy (anonymous):

1+sec^2 x sin^2 x=sec^2 x =1+(1/cos^2 x) sin^2 x = 1/cos^2x =cos^2 x/cos^2 x + sin^2 x/cos^2 x = 1/cos^2 x = (cos^2x+sin^2 x)/cos^2 x = 1/cos^2 x

OpenStudy (anonymous):

what was your problem to begin with ?

OpenStudy (anonymous):

1 + sec2x sin2x = sec2x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (anonymous):

1 + sec²x sin²x = sec²x 1 + (1/cos²x) sin²x = sec²x 1 + (sin²x/cos²x) = sec²x 1 + tan²x = sec²x verified.

OpenStudy (anonymous):

Or if you don't consider 1 + tan²x = sec²x to be an identity, then 1 + tan²x = sec²x (cos²x/cos²x) + (sin²x/cos²x) = sec²x (cos²x+sin²x)/cos²x = sec²x and knowing that `cos²x+sin²x=1`, 1/cos²x = sec²x sec²x=sec²x

OpenStudy (anonymous):

Will see what Lars is going to say about this :)

OpenStudy (larseighner):

\[ \large \begin{align} 1 + \sec^2x\sin^2x &= \sec^2x \cr &= 1 +\tan^2x \cr &=1+ {\sin^2x \over \cos^2x} \cr &= 1 + {1 \over \cos^2x}\sin^2x \cr &=1 + \sec^2x\sin^2x\end{align} \]

OpenStudy (anonymous):

Yes, the same thing, but working the other side.

OpenStudy (anonymous):

Just seems a bit wild though, to start verifying with making the simpler side more complicated rather than simplifying the hard side, but I like this approach to .

OpenStudy (larseighner):

What is important in demonstrations like this is to work only one side. Otherwise you end up with 1=1 or 0=0 which is not a very good demonstration unless you turn it upside down and start with 1=1.

OpenStudy (larseighner):

The reason I started on the right side was that this Pythagorean identity has been tattooed on my brain (and it is also one that exercise authors think is obscure, so they use it a lot).

OpenStudy (anonymous):

Okay... I see your approach just the same way. It's probably just that I am thinking in a standard way and you are very creative!

OpenStudy (anonymous):

Thank you guys... I'm still in the process of understanding each step.

OpenStudy (anonymous):

Good luck:)

OpenStudy (anonymous):

Thanks. :D

OpenStudy (larseighner):

Ok, 1+tan^2 = sec^2 is a Pythagorean Identity. Here is a geometric interpretation. See the pink tangent and the secant is cos + versine + exsecant on the horizontal line.

OpenStudy (larseighner):

The rest of the lines are basic trig identities.

OpenStudy (larseighner):

Back soon. Have to reboot. Something (I suspect Opera) is leaking memory like a sieve.